# Modeling of Far-Field Quantum Coherence by Dielectric Bodies Based on the Volume Integral Equation Method

### v2

@@ Abstract
Keywords Hong-Ou-Mandel (HOM) effect $\cdot$ two-photon interference $\cdot$ volume integral equation (VIE) $\cdot$ far-field scattering $\cdot$ second-order correlation function $\cdot$ coincidence counts
1 Introduction
@@ 2 Theory
2.1 The correlation function in the frequency domain
@@
and similarly for $\hat{E}^{(+)}_{b,\mathrm{NA}}(t)$, where $W(\mathbf{k})$ is the angular sensitivity function of the detector. The experimentally relevant quantities are then $G^{(1)}_{a,\mathrm{NA}}=\langle\hat{E}^{(-)}_{a,\mathrm{NA}}\hat{E}^{(+)}_{a,\mathrm{NA}}\rangle$, $G^{(2)}_{ab,\mathrm{NA}}=\langle\hat{E}^{(-)}_{a,\mathrm{NA}}\hat{E}^{(-)}_{b,\mathrm{NA}}\hat{E}^{(+)}_{b,\mathrm{NA}}\hat{E}^{(+)}_{a,\mathrm{NA}}\rangle$, followed by $g^{(2)}_{ab,\mathrm{NA}}=G^{(2)}_{ab,\mathrm{NA}}/(G^{(1)}_{a,\mathrm{NA}}G^{(1)}_{b,\mathrm{NA}})$.
Substituting (1213) into (9), one obtains
@@
2.2 The normalized coincidence number function in the time domain
@@
2.2.1 Beam-splitter reference formulation
@@
2.2.2 Extension to an arbitrary lossless scattering environment
@@
2.2.3 Normalized coincidence number
@@
so that the dip width is governed by the coherence time $\sigma$.
The temporal envelope of the excitation pulses thus defines the coherence time of the generated photons and determines the characteristic width of the interference curve. In this work, we emphasize the comparative behavior of $\tilde{N}_{c}(\delta\tau)$ across different scattering environments; the absolute temporal scale can be mapped to a specific experimental implementation through the pulse bandwidth and detector timing response.
2.3 Numerical method
@@ 3 Results
In this section, two representative scenarios are investigated: (i) two-photon scattering from a dielectric sphere, which exhibits angle-dependent HOM-type suppression of coincidences; and (ii) the generation of spatial–polarization entanglement mediated by a space-variant Pancharatnam–Berry phase metasurface. In both cases, the structure acts as a distributed three-dimensional beam splitter: the effective “path amplitudes” are encoded in the continuous angular dependence of the far-field response rather than in discrete optical paths.
3.1 Dielectric Sphere
@@
3.1.1 Validation against analytical reference
@@
3.1.2 Classical fourth-order field correlation for comparison
@@
3.1.3 Time-domain coincidence curves
@@
Figure 4: (a) Four representative detection configurations on the angular map. (b) Normalized coincidence counts $\tilde{N}_{c}$ versus the controlled input delay $\delta\tau$ for these configurations.
3.2 Dielectric Metasurface
@@
Figure 7: Normalized coincidence counts $\tilde{N}_{c}$ between the two detection positions versus the controlled input delay $\delta\tau$.
4 Conclusion

### v3

@@ Abstract
I Introduction
@@ II Theory
II-A The correlation function in the frequency domain
@@
and similarly for $\hat{E}^{(+)}_{b,\mathrm{NA}}(t)$, where $W(\mathbf{k})$ is the angular sensitivity function of the detector. The experimentally relevant quantities are then $G^{(1)}_{a,\mathrm{NA}}=\langle\hat{E}^{(-)}_{a,\mathrm{NA}}\hat{E}^{(+)}_{a,\mathrm{NA}}\rangle$, $G^{(2)}_{ab,\mathrm{NA}}=\langle\hat{E}^{(-)}_{a,\mathrm{NA}}\hat{E}^{(-)}_{b,\mathrm{NA}}\hat{E}^{(+)}_{b,\mathrm{NA}}\hat{E}^{(+)}_{a,\mathrm{NA}}\rangle$, followed by $g^{(2)}_{ab,\mathrm{NA}}=G^{(2)}_{ab,\mathrm{NA}}/(G^{(1)}_{a,\mathrm{NA}}G^{(1)}_{b,\mathrm{NA}})$. Note that in our numerical implementation, under the narrowband approximation, the spatial integration over the numerical aperture is not explicitly performed. Instead, by approximating the collected signal through contributions from specific discrete modes at the center of the detection angles, the substitution of (1213) into (9) yeilds:
@@
II-B The normalized coincidence number function in the time domain
@@
II-B1 Beam-splitter reference formulation
@@
II-B2 Extension to an arbitrary lossless scattering environment
@@
II-B3 Normalized coincidence number
@@
so that the dip width is governed by the coherence time, represented by the Gaussian width parameter $\sigma$. It should be explicitly clarified that $\sigma$ is an assumed constant, independent of the monochromatic frequency-domain scattering results. This parameter is employed solely to enable the transition from monochromatic data to a temporal profile; thus, the resulting coincidence curves do not represent a full broadband frequency-domain simulation, but rather a theoretical visualization based on the assumed temporal envelope.
The temporal envelope of the excitation pulses thus defines the coherence time of the generated photons and determines the characteristic width of the interference curve. In this work, we emphasize the comparative behavior of $\tilde{N}_{c}(\delta\tau)$ across different scattering environments; the absolute temporal scale can be mapped to a specific experimental implementation through the pulse bandwidth and detector timing response. Accordingly, $\sigma$ is assigned as a dimensionless value consistent with this normalized temporal scale.
II-C Numerical method
@@ III Results
In this section, two representative scenarios are investigated: (i) two-photon scattering from a dielectric sphere, which exhibits angle-dependent HOM-type suppression of coincidences; and (ii) the generation of spatial–polarization entanglement mediated by a space-variant Pancharatnam–Berry phase metasurface. In both cases, the structure acts as a distributed three-dimensional beam splitter: the effective “path amplitudes” are encoded in the continuous angular dependence of the far-field response rather than in discrete optical paths. Specifically, $\sigma=1.2$ is adopted to facilitate the temporal visualization of the HOM dip. This choice ensures a clear representation of the coincidence features within the chosen time window, independent of the frequency-domain scattering calculations.
III-A Dielectric Sphere
@@
III-A1 Validation against analytical reference
@@
III-A2 Classical fourth-order field correlation for comparison
@@
III-A3 Time-domain coincidence curves
@@
Figure 4: (a) Four representative detection configurations on the angular map. (b) Normalized coincidence counts $\tilde{N}_{c}$ versus the controlled input delay $\delta\tau$ for these configurations. The Gaussian width parameter used for the temporal profile is $\sigma=1.2$.
III-B Dielectric Metasurface
@@
Figure 7: Normalized coincidence counts $\tilde{N}_{c}$ between the two detection positions versus the controlled input delay $\delta\tau$. The Gaussian width parameter used for the temporal profile is $\sigma=1.2$.
IV Conclusion